OpenStudy (anonymous):

−6x+ y =20 −3x−3y =−18 can anybody explain how to do substitution?

jimthompson5910 (jim_thompson5910):

solve the first equation for y −6x+ y =20 y =20 + 6x Then plug this into the second equation −3x−3y =−18 −3x−3(20+6x) =−18 -3x - 60 - 18x = -18 then solve for x

jimthompson5910 (jim_thompson5910):

use the solution for x to find the solution for y

OpenStudy (anonymous):

but shouldnt i end up with an answer for X & Y?

jimthompson5910 (jim_thompson5910):

yes, what did you get when you solved -3x - 60 - 18x = -18 for x

OpenStudy (anonymous):

-3, is that correct? & then I plug in that to either one of the equations to get y?

jimthompson5910 (jim_thompson5910):

no that's not correct

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

-3x - 60 - 18x = -18 -21x - 60 = -18 -21x = -18+60 -21x = 42 x = 42/(-21) x = -2

OpenStudy (anonymous):

ohhhhh, i see!

jimthompson5910 (jim_thompson5910):

now use this to find y

OpenStudy (anonymous):

thank you!